OpenStudy (anonymous):

PLEASE HELP?! The equation of the circle with center (3, -2) and radius 7 is:

OpenStudy (anonymous):

\[(x-h)^2+(y-k)^2=r^2\] replace \(h\) by \(3\), replace \(k\) by \(-2\) and replace \(r\) by \(7\)

OpenStudy (anonymous):

(x-3)^2 + (y+2) = 7^2

OpenStudy (anonymous):

looks good to me

OpenStudy (anonymous):

well, you forgot the square on the second term

OpenStudy (anonymous):

right... just noticed that but thanks so much :) &^2 = 49

OpenStudy (anonymous):

\[(x-3)^2 + (y+2)^2= 49\]

OpenStudy (anonymous):

yup